OpenStudy (lilai3):

@whpalmer4 Factor the polynomial. -81 + a^4 Thank you! I will reward by giving you a medal! ^_^

OpenStudy (whpalmer4):

Hey, wasn't I going to get an Oreo for the last one? :-)

OpenStudy (lilai3):

AH this one seems too confusing...

OpenStudy (whpalmer4):

Okay, \[-81+a^4\]Any ideas? How about we rewrite it in the other order, just to see if that makes anything jump out at us...

OpenStudy (lilai3):

hmm? oh yes oreo... okay, what's your address? i'll send it to you! (: jk i don't wanna know..

OpenStudy (whpalmer4):

Eat it for me, and enjoy it :-)

OpenStudy (lilai3):

the greatest common factor that i can only think of is 1 right now.

OpenStudy (whpalmer4):

did you try rearranging it as I suggested?

OpenStudy (lilai3):

lol thanks, palmer. you are making my day! (:

OpenStudy (lilai3):

a^4-81?

OpenStudy (lilai3):

Nothing "jumped out" at me...

OpenStudy (whpalmer4):

Yeah. Does that give you any ideas?

OpenStudy (whpalmer4):

What's the square root of 81?

OpenStudy (lilai3):

9, sir.

OpenStudy (whpalmer4):

What's the square root of \(a^4\)?

OpenStudy (lilai3):

...

OpenStudy (lilai3):

i don't know...a^2?

OpenStudy (whpalmer4):

\[\sqrt{a*a*a*a} = a*a\] right?

OpenStudy (whpalmer4):

my cat says "if you're not getting an Oreo, you can feed me now!" brb

OpenStudy (lilai3):

lol <3 aww you have a kitty :3

OpenStudy (lilai3):

so the square rot of a^4 is a^2 right? if it is, then in that case, me correcto

OpenStudy (whpalmer4):

right, if \(a>0, \text{ then } \sqrt{a^4} = a^2\)

OpenStudy (lilai3):

I see. What do I do now?

OpenStudy (whpalmer4):

Well, remember the difference of squares? \[(a^2-b^2) = (a+b)(a-b)\] We have \[a^4-81 = (a^2)^2 - (9)^2\]

OpenStudy (whpalmer4):

maybe it would be less confusing if I said the difference of squares pattern is \[(x^2-y^2) = (x+y)(x-y)\]

OpenStudy (lilai3):

OHHH. I see now.!!!

OpenStudy (lilai3):

wait now i am totally confused of what to do next

OpenStudy (whpalmer4):

Well, \[(x^2-y^2) = (x+y)(x-y)\]We have \[(a^4-81) = ((a^2)^2-(9)^2)\]So \(x=a^2,\,y=9\)

OpenStudy (lilai3):

it looks so confusing, but somehow i get it

OpenStudy (whpalmer4):

so write it out, what does that give us?

OpenStudy (lilai3):

(a^2-9) (a^2 +9) ?

OpenStudy (lilai3):

is that right?

OpenStudy (whpalmer4):

Good so far. Are we done?

OpenStudy (whpalmer4):

I guess I kind of let the cat out of the bag by saying "so far" :-)

OpenStudy (lilai3):

what do you mean? yes i think so...with this problem, at least

OpenStudy (whpalmer4):

Look again...

OpenStudy (lilai3):

okay then i do not understand..

OpenStudy (whpalmer4):

Okay, can you factor either of \((a^2+9)\) or \((a^2-9)\)?

OpenStudy (lilai3):

..no...?

OpenStudy (whpalmer4):

Ask grandpa :-)

OpenStudy (lilai3):

he's asleep

OpenStudy (whpalmer4):

What's 1*1?

OpenStudy (whpalmer4):

What's 2*2?

OpenStudy (whpalmer4):

What's 3*3?

OpenStudy (lilai3):

1....if my memory is correct...which i am hoping it is.... 2x2=4 9

OpenStudy (whpalmer4):

What's a*a?

OpenStudy (whpalmer4):

I might have to ask for my Oreo back :-)

OpenStudy (lilai3):

a^2

OpenStudy (whpalmer4):

My cat just said "mrrrow"...which is cat-lish for "isn't that a difference of squares?"

OpenStudy (whpalmer4):

\[(a^2-9) = ((a)^2-(3)^2)\]

OpenStudy (lilai3):

don't you have to add somehting because it is part o that formula thingy?

OpenStudy (whpalmer4):

\[x = a, y = 3\]\[(x^2-y^2) = (x+y)(x-y)\]

OpenStudy (lilai3):

yes, that so.. [(a-3) (a-3)]^2?

OpenStudy (whpalmer4):

No :-( \[x^2-y^2 = (x+y)(x-y)\]\[x=a,\,y=3\]\[x^2-y^2=(x+y)(x-y) = (a+3)(a-3)\]

OpenStudy (whpalmer4):

We can't factor \((a^2+9)\) because it is a sum of squares, not a difference. \[a^4-81 = (a^2+9)(a^2-9) = (a^2+9)(a-3)(a+3)\]

OpenStudy (lilai3):

this problem is confusing me

OpenStudy (whpalmer4):

Any time you see something squared, and a minus sign, you need to think difference of squares?

OpenStudy (lilai3):

yes

OpenStudy (whpalmer4):

Let's multiply that expression out and see what we get: \[(a^2+9)(a-3)(a+3) = (a^2+9)(a^2 - 3a + 3a -3*3) = (a^2+9)(a^2-9) \]\[=a^2*a^2 + 9a^2 - 9a^2 -9*9\]\[ = a^4 - 81\]

OpenStudy (lilai3):

ohhh

OpenStudy (lilai3):

thanks!

OpenStudy (whpalmer4):

We can't factor the \(a^2+9\) because if we have \[(x+c)(x+d)\]we get those pesky +x*d + x*c and they only cancel if one of them is negative, and they have the same absolute value

OpenStudy (whpalmer4):

Even \[(x+c)(x+c)\] doesn't work because\[(x+c)(x+c) = x^2 + cx + cx + c^2\]Need one of those to be \(-cx\) but if that happens then we have a \(-c^2\) at the end. We can actually do it if we are willing to introduce imaginary numbers into the game. Have you heard of them yet?

OpenStudy (lilai3):

uh no

OpenStudy (usukidoll):

perfect square and there's a part of it that you need to factor again

OpenStudy (usukidoll):